UT Nov 28 2021 - Split night PEPSI/PETS w/DEC



Night Info

  • Sunset 17:19
  • Twlight Ends 18:40
  • Handover to PEPSI 20:15
  • Moon rises 00:35
  • Handover from PEPSI 02:07
  • Twilght Starts 05:35

Night Plan

AO

  • Configure for LUCI-Bino, authorize LUCI-1
    • Location of AO/ESM scripts: /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM
    • AO OT file: AO_Testing_V37.xml
    • ESM OT file: ESM/ESM_V27.xml
      • mixed mode scripts have already been generated and corrected to include SL on LUCI-2 - exported to sub-folders based on target name
    • Rotator Info - To translate between Instrument Rotator and PA
      • ROTANGLE = POSANGLE - PARANGLE + (180 - LEFTZEROPOINT)
      • Lett ZP =-187.735
  • LUCI-1 AO Tests start at twilight until 20:15
    • Start with OSCO in twilight
      • First pass without NCPA
      • Second pass with NCPA
    • Close the loop on Bin 1 & Bin 2 stars (Bin 4 still not available)
      • RA 23hr-0hr Folder in OT File
      • Scripts execute in-field dithers (1") then out of patrol field
      • Scripts have 60 deg PAs to account for bad rotator angles
      • 3xBin1 stars, 2xBin2 stars, 2xBin4 stars
      • Science Target:
        • Bin 4 - Einstein's Cross - Ks -15 dithers - PAs available every 60 deg
          • generate from OT - ANG/ULIRGs/LIRGs Folder/Einsteins Cross NEW
          • Option to select the best PA/Rotator range
        • HR 8799 ADI test - use pre-generated scripts - cannot generate from OT file without Beta Version of OT
          • Files; /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM/AO/ADI_Tests
          • run collimation script first
          • Science script needs 256x2048 sub-framing
  • If OSCO tests show seeing too poor to continue -AND >30min before PEPSI/PETS handover execute backup LUCI science
    • File Location: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers/COBandhead_Mergers_V31.xml
    • NGC 6598 - observe telluric first
      • G210 CWL 2.3microns 0.5" slit - 300sec exposures
  • Reconfigure from PEPSI/PETS to bino-LUCI w/LUCI-1 authorized- 01:37-01:52hr
  • After handover from PEPSI continue with LUCI-1 AO testing, ESM testing
    • Run another OSCO to estimate seeing and system performance
      • First pass without NCPA
      • Second pass with NCPA
    • Bin 1 & Bin 2 Close the loop on stars. Bin4 can be tested as well.
      • RA 04-05hr Folder in OT
      • RA 08-09hr Folder in OT
      • Mrk 231- on-axis science target Bin2 (R=13.3)
        • visible 03:20-twilight
        • lower priority - execure if time available
        • BrGamma to avoid saturating the central pixels and still use full FOV
        • 17 dithers - 12 on, 5 off
        • PA = 45 - can probably regenerate with slighlty different PAs on-the-fly if OCAM rails
        • PSF star to be taken after target (9 dithes in-field -WD of similar flux)
          • Future ADI target - transit occurs after sunrise
    • ESM Testing (Bin 1 or 2 ESM)
      • HCG 31 (R=10.7 AO Ref star Bin 2 PA=0) - target hits 30deg elevation limit at 04:15hr
        • Script cycles through KHJ filters, 10 dithers (6on 4off) - execute as much as needed
      • M67 (R=6.5 AO Ref Star Bin 1) - 00:30-twilight
        • Script is Br Gamma (to avoid too many saturated stars - AO Ref will be saturated) - 11 dithers - 7 on 4 off
        • Scripts for PA=0,60,120,180,240,300
        • ~15 min per PA
      • If Bin 4 becomes available
        • Arp 143 (Bin 4 forced to R=16.51)
          • Cycles through KHJ - 22:30-twilight
          • Total time ~50min
        • M67 (Bin 4 forced to R=16.51)
          • Cycles through KHJ (10dithers 6 on 4 off)
          • Total time ~50min
  • If conditions are too poor for AO/ESM (after OSCO afttempt) 02:15hr-onwards
  • Continue to stay with LUCI-Bino in the beam to monitor seeing, if things improve, switch back to AO
    • CaTriplet in QSOs - 02:15hr until NGC 4004 or Dwarf Merger rises above 30deg elevation - OR - AO conditions
      • /home/lbto/Service_Observing/2021B_D_E/PoorWeather/QSOs/RLRQQSO_V4.xml
      • Any conditiosn
      • HS_0624+6907
        • G210 0.5" slit J-band CWL 1.17 LUCI-bino PA=90
        • Script is 30min shutter time - repeat as weather deteriorates
        • Run script 2x if mono-LUCI
        • Visisble 20hr-twilight (Dec +69)
        • Run telluric first
    • Local Mergers
      • /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers/COBandhead_Mergers_V31.xml
      • any conditions
      • NGC 4004 - visible 03:30hr-twilight
        • Brightest point in center should be nucleus (there will be other clumps - SF regions)
        • script is 30min shutter time Bino - increase repeates depending on conditions
        • G210 0.5" slit CWL 2.27microns
    • Dwarf Merger
      • /home/lbto/Service_Observing/2021B_D_E/LUCI/Rothberg_Science/Dwarf_Mergers/DwarfNergers_OT_V4.xml
      • SDSSJ1148+2149
        • up to ~1mag extinction, seeing up to 2"
        • Fraternal L1 zJsepc, L2 HKspec - if one LUCI is offline, then observer HKspec only
        • Telluric star first (rises before target)
        • Target rises above 30deg elevation 03:15hr
        • Run script 2x (script is 30min shutter time more if time permits)
    • Photometric Standards
      • Any seeing - sky must be clear/photometric
      • OT File: /home/lbto/Service_Observing/2021B_D_E/LUCI/LUCI_2021_Restart_V1.63.xml
        • Pick stars visible - try to sample at least 2 airmasses
        • Script runs through EVERY broadband filter - this will take time with dithers

Non-AO

In case AO has technical issues (non-weather related) at start of night
  • Configure for bino-MODS
  • Feige 110 SpecStd in twilight
  • M31 Nova until 20:15hr
  • After PETS - Reconfigure for...TMS testing

Night Log

Preparations

23:40 Check FS alignment on LUCI1 in preparation for OSCO.

luci1.20211128.0301.fits - N30, N30_FS, clear+KS. The FS is not far off, but it is not aligned with gaps at top and right.

luci1.20211128.0302.fits - N3.75, N30_FS. FS centered at 1021.5, 1018.75

Went back to N30 - luci1.20211128.0303.fits with N30+N30_FS used to align. 0304 is confirmation.

For kicks, where is FS on N3.75? image 305 has FS centered @ 1027.25, 1031.5. Pretty close to the center.

23:58 Done with FS alignment. Doug and Steve are prepping AO system.

Log

00:20 Open -

00:36 Steve is checking pointing and collimation on both LUCIs in the event that AO cannot be done and we back-up to binoLUCI for science.

Note: Doug is keep the AO log for 20211127_SX here https://wiki.lbto.org/AdaptiveOptics/20211127_SX_Night

OSCO

00:47 Sending the OSCO script. CCD47 view looks odd on Doug's display - but OK for Steve on obs1.

00:48 NCPA off for first set of data
NCPA OFF L1  
  307-316,317-326 offset bt 316&317

01:01 NCPA on
NCPA ON L1   sent detxy abs -5 5 manually
  327-335,336-345   FWHM 2.8 pix, 0.04"

Juan Carlos reported SR 55% for star

Bin 1 test

01:13 Slewing to Star 2 R=11.1 (catalog; 12.1 predicted for OCAM) H=9.1, PA=60 (AO_ESM/AO_Testing_V37.xml) --- 01:17 we're at the target, now starting to close the loop.
  • Star 2 is G216-35 (an M0 star). OCAM says R=12.6
  • 01:23 Turning on NCPA - starting to take data.
    • 4.4 x 3.56 pix for image 5347.
    • Dithering with 1" offsets first, then off.
  • TN 013458
  • 5351
  • 01:36 -- between 5352 and 5351 we are making a 100",100" offset so now we are out of both the AO & Active Optics patrol field.
  • 01:39 --- between 5353 and 5352 we moved back to 0,0 offset and loop closed (all went well).
  • TN 013943
  • 01:42 --- just returned from a big offset and again, the loop closed as expected.
  • 01:44 zen corrected DIMM is ~0.9" (elev=80 deg).
  • Doug "after the large offset, the camera lens values are large." He says it's not causing a problem now, but it means that something is moving when we make the offset that needs to be investigated.
  • TN 015438
  • L1 images 346 - 361
  • Juan Carlos reports - last image has SR 45%

01:57 Slewing to Star 2 but now with PA=0 so that the Active Optics GS will remain in its patrol field after the big offset.
  • 02:01:42 Optimize gain failed with safe skip, then succeeded.
  • Safe skipped right afterwards
    • Weird optimize gain plot...
    • OCAM says R=12.6 again.
  • 02:05 Done, NCPA is applied and we are taking data, starting with image 362.
    • Image 5362 shows a jump. Barry says there is a 40-pix jump around frame 8 in the cube. Time 2021-11-28T02:04:56.5213 + 2.51*8
  • 02:17 Large offset (between 367 & 368). Active optics has taken over for guiding and WFSing.
  • 02:19 Came back - the guide star was close enough that the loop re-closed without a problem. But the new Pause/Resume closes loop first with 10-mode gain and then adding the HO gains, so this helps the problems we used to have with a bunch of skip frames after coming back from an offset. (Note - see below ~11:15 UT - this version wasn't yet implemented, we were just lucky).
  • L1 images 362 - 377
  • FWHM on Aladin stack 5377 is 3.4 pix

Bin 2 test

02:38 Slewing to Star 6 R=12.8 (catalog; 13.8 predicted for OCAM) H=9.4, PA=0
  • Star 6 = G71-13 (M4 dwarf). Steve says it looks like a double (fairly close & same mag)

02:45 Slewing to Star 5 R=13.3 (catalog; 14.3 predicted for OCAM) H=9.7, PA=0
  • Star 5 = G172-41 (M3 star).
  • 02:53 - optical gain has a very small value (0.1 vs 0.3), a factor of 3 lower than with bin 1
  • OCAM says 13.6-13.7
  • Doug is noting in the log the elongated, round images.
  • L1 images 378 - 389.
  • FWHM on 5389 is 4.6 x 3.6 pix.
  • Camera lens is all the way to one end. Just barely out of the lens range - did it start this way or move closer to the edge as the field rotated?
  • LFBG is at 150 deg right now. (it was 136 when we started).

03:20 Reconfiguring for PEPSI

08:35 Reconfiguring back to the LUCIs

There was GCS problem - corrected @ 08:53

OSCO
08:54 Running OSCO script. R=7.46 catalog; OSCO says 8.6 * DIMM Seeing 0.82" (zen)
NCPA status L1 comments
NCPA off 390-409 01:58 skip frames
  ~399 Pupils look like "spinning wheels" - overdriving gains

After script ended, Steve optimize the gain again to if that would help with the 'overdriving' problem.

Doug is making a test to see if there is a rotational offset. The reconstructor is measured at a certain angle of the re-rotator and the WFS data have to be referenced to this angle. Maybe we are not applying the correct angle? (there may be a temperature dependence). By making an offset in the re-rotator by 1 deg, the "spinning wheels" went away.

09:25 Doug made a change to a file. Steve is repeating the NCPA-off OSCO. adSec failed. Doug is taking care of the recovery.

09:43 (just before) - Resent the OSCO script, but there is a problem - Canceling and resending preset.

09:59 No commands are being received to move the camera lens (because stop/start WFS arbitrator?). Canceling preset and Doug is restarting WFS arbitrator, and all WFS processes.

10:08 Resending preset. Camera lens is moving. During optimize gain, a skip frame (10:13) and "spinning wheels" are worse now than before. It seems like the value that was there before is the best.

10:19 Steve canceling preset. Resending. Finally taking data, but decided to take the data with NCPA.

NCPA on 410-429   DIMM 0.83" (zen)
  422 3.12 x 3.01 pix

Bin 2 test (offsets keep off-axis star within off-axis patrol field)
10:38 (before) - Sending preset to Star 5 from 08-09hr set R=12.8 (catalog; 13.8 predicted for OCAM) H=9.7, PA=0
  • OCAM says 13.0 (M3 star) G117-56
  • subwindowed to 256 x 2048: subwin 1 1052 1 256 2048. DIT = 0.31. ~5700 ADU in the peak.
  • 03:54 resumed script. DIT x NDIT = 0.31 x 100 LIR, cube
  • NCPA is on. Image 5432 has some evidence of trefoil. LFBG rotator at 88 deg.
  • 04:02 offset and back worked fine.
  • L1 430-440
  • 440 - loop opened on return from sky offset to target. Just before 11:10 UT. Doug says that we are not using the "ramp or step gain" now - that's in the new version which Brandon sent us instructions on how to start. Rather than re-closing the loop, we are going to try the new version.

11:29 Resending Star 5 (08-09hr) PA=0 script now with new SW (we thought we were, but later realized it is not doing the "step gain" after an offset.
  • During the first "acquisition" image, 441, the star jumped by 0.15".
  • subwin 1 1031 1 256 2048
  • 11:45 running script
  • 11:50 - between 447 and 448 return from large offset. Skip frame.
  • 11:57 - Restarting from an "on"-position to watch the behavior when returning from a large offset. This time it was successful, but there was not so much tip this time. The "step gain" does not seem happening.
  • 12:04 - returning from a large offset, on image 453 the loop opened.
  • 12:06 - restarting from an "on" position. The loop opened again upon returning.
  • 12:15 - trying again from the "on" position. This time, In the IIF Controls, Steve set the "Ignore AO Rip" to "Pause GCS" in an attempt to stop guiding in the offset position. GCS is not sending centroids - issued a warning - but then started guiding and WFSing. Nevertheless, this time when we got back on source, the loop closed.
  • 12:20 - Steve changed to "Notice RIP". Again this caused a warning after which guiding and WFS were immediately reenabled.
  • L1 441-461

As a result of these problems when active optics passes off guiding/collimation to AO, we need to create scripts with offsets that, if they move the AO ref star out of the AO patrol field, also move the off-axis star out of its patrol field.

Bin 4 ESM+SL test
12:25 Steve is reauthorizing for both LUCIs so we can run the next ESM+SL script.

M67_PA180_Bin4_ESM_SL.xml

The catalog AOref star magnitude is 16.51 (catalog mag is actually 16.35), but OCAM says 17.4.

Camera lens is pegged at 100. Rotator at 157. Doug thinks that the main problem in this case is the faintness of the star, but we will try a new PA.

M67_PA270_Bin4_ESM_SL.xml

12:47 Slewing to M67 PA=270. Now the camera lens is way off of the limit. The LFBG rotator is at 242 deg. Trying to optimize gain - Doug will do this manually - 0.6 and 0.25.

12:57 Skip frame. Loop opened. Camera lens railed at 100 again. The pupil is so dim that the WFS gets bogus measurements and pushes the lens off to the edge.

The AO ref star was just too faint. Bin 4 ESM may have worked with a brighter star, though.

Bin 1 ESM+SL test

M67_Bin1_PA0_ESM_SL.xml

13:00 Sending the preset to M67, PA=0 and using an R=8.96 (catalog) A0 ref star. OCAM says 10.3.

13:19 Loop is closed and gains have been optimized. Starting to collect data.

L1 462-

13:23 Jump on image 462 to the NE by about 0.4".

After this, the off-axis guide star jumped out of the field (similar to what we saw earlier with PEPSI).

13:24 Large offset request made (-330, 390) and copointing limit exceeded. Doug opened loop.

13:27 Sending the preset again. FWHM on DX guider is 0.7" (DIMM 0.8") Doug is working on closing the loop - but it is getting pretty bright.

13:36 Closed the loop. Loop kept trying to rail the camera lens (LFBG at 321 deg) though. Doug things that the high background introduced some shifts.

13:39 We lost the off-axis guide star again, but the AO ref star is still there and the loop is closed. Off-axis guiding continues on the right.

Topic revision: r10 - 28 Nov 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback